OpenStudy (anonymous):

Let f(x) = –4x + 7 and g(x) = 10x – 6, find f(g(x)). –40x + 70 –40x + 64 –40x + 24 –40x + 31

OpenStudy (anonymous):

You want to find f(g(x)). So you simply start by placing g(x) everywhere where there is a "x" in f(x) function, so: \[f(g(x))=-4(10x-6)+7\] Then you simplify

OpenStudy (saranya):

g(x) = 10x - 6 f(x) = -4x + 7 f(g(x)) = -4(10x - 6) + 7 f(g(x)) = -40x + 24 + 7 f(g(x)) = -40x + 31

OpenStudy (anonymous):

thanks